# Neuron Empirical Gradient: Connecting Neurons’ Linear Controllability and Representational Capacity

### v1

@@ E.2 More Data about Efficiency
We report the accuracies of major-vote probers with different neuron sizes for all tasks to provide additional evidence for the discussion about the representation and acquisition efficiency of skill neurons in § 6.1. The results are demonstrated in Figure 13 and Figure 14 for zero-shot and few-shot prompting settings.
@@
E.3 Probing With Varying Neuron Sets
We report the aggregated accuracies across all 22 tasks in MCEval8K in Figure 15 to provide additional evidence for discussion in § 6.3. It demonstrates that many neurons can construct the classifiers in solving the language tasks, showing their ability to represent language skills and knowledge.
E.4 Tree-prober: Flatteness vs. Hierarchy
To investigate the balance between hierarchy and independence of skill neurons, we train Tree-prober with fixed neuron features ($2^{10}$) but with different depths and trees. For each task, we train 10 Tree-probers, varying the number of trees (#n_tree $\in(2^{0}\sim 2^{9})$) and the tree depth (#n_layer $\in(2^{10}\sim 2^{1})$), which fewer trees with deeper layers indicate a more hierarchical structure. All tasks show a camel curve given stronger hierarchies. We report the optimal #n_layer for different tasks as follows: CSQA(4), MNLI(16), SWAG(16), Stereoset(16), Agnews(32), MyriadLAMA(32), mLAMA(32), XNLI(32), POS(64), FEVER(64), Toxic(64), LTI(64), GED(128), IMDB(128), M-Amazon(128), CHUNK(256), NER(256), Amazon(256), PAWS(256), HaluEval(256), M-POS(256), TempLAMA(1024). This demonstrates that Different language skills require different hierarchy levels. For instance, factual tasks benefit from flatter structures, while linguistic tasks prefer deeper hierarchies.
For all tasks in MCEval8K, we plot the accuracies of trained models with varying hyperparameters, including the number of trees and layers per tree. The number of trees is set to $2^{N}$, where $N$ ranges from 0 to 10, and the number of layers is set to $2^{M}$, where $M$ ranges from 1 to 11. Training is conducted only for configurations where $N+M<12$. The results are visualized as 3D surfaces, where the x-axis represents the logarithm of the number of trees (#log_ntree), the y-axis shows the logarithm of the number of layers (#log_nlayer), and the z-axis indicates the accuracy evaluated on the test set. We display the results for all tasks under the zero-shot setting in Figure 17 and those under the few-shot setting in Figure 18.
Figure 17: Accuraries of trained random forest models with the zero-shot setting on Llama2-7B.
Figure 18: Accuraries of trained random forest models with the few-shot setting on Llama2-7B.
Appendix F Prompting Setups
In this section, we list all the instructions we use for each task in MCEval8K. It includes design instructions, options, and a selection of few-shot examples. As mentioned in § 5.2, we adopt two instruction settings, zero-shot and few-shot. For few-shot prompting, we set the number of examples to the same number as the number of options and ensure each option only appears once to prevent majority label bias Zhao et al. (2021). All the few-shot examples are sampled from the training set. Finally, we list all the instructions and options we used for skill neuron probing examples by showing one zero-shot prompt.
@@ Appendix F Prompting Setups
Appendix G Diverse Contexts for Skill Neuron Generality Evaluation

### v2

@@ I.2 More Data about Efficiency
We report the accuracies of major-vote probers with different neuron sizes for all tasks to provide additional evidence for the discussion about the representation and acquisition efficiency of skill neurons in § H.1. The results are demonstrated in Figure 16 and Figure 17 for zero-shot and few-shot prompting settings.
@@
I.3 Probing With Varying Neuron Sets
We report the aggregated accuracies across all 22 tasks in MCEval8K in Figure 18 to provide additional evidence for discussion in § H.3. It demonstrates that many neurons can construct the classifiers in solving the language tasks, showing their ability to represent language skills and knowledge.
Appendix J Prompting Setups
In this subsection, we list all the instructions we use for each task in MCEval8K. It includes design instructions, options, and a selection of few-shot examples. As mentioned in § 7.2, we adopt two instruction settings, zero-shot and few-shot. For few-shot prompting, we set the number of examples to the same number as the number of options and ensure each option only appears once to prevent majority label bias Zhao et al. (2021). All the few-shot examples are sampled from the training set. Finally, we list all the instructions and options we used for skill neuron probing examples by showing one zero-shot prompt.
@@ I.3 Probing With Varying Neuron Sets
Appendix K Diverse Contexts for Skill Neuron Generality Evaluation